Constants

The following values may be useful when solving this tutorial.
Constant Value
E∘Cu 0.337 V
E∘Zn -0.763 V
R 8.314 J⋅mol−1⋅K−1
F 96,485 C/mol
T 298 K

Part A

Part complete

In the activity, click on the E∘cell and Keq quantities to observe how they are related. Use this relation to calculate Keq for the following redox reaction that occurs in an electrochemical cell having two electrodes: a cathode and an anode. The two half-reactions that occur in the cell are

Cu2+(aq)+2e−→Cu(s)

and

Zn(s)→Zn2+(aq)+2e−

The net reaction is

Cu2+(aq)+Zn(s)→Cu(s)+Zn2+(aq)

Use the given standard reduction potentials in your calculation as appropriate.

Express your answer numerically to three significant figures.

View Available Hint(s)

--

SubmitPrevious Answers

Keq

K e q

=
Answer #1

cell = E°cathode - E°anode = 0.337V -(-0.763V)

= 1.100V

= 1.10V

Ecell = E°cell - RT/nF × ln(Keq) = 0

cell = RT/nF × ln(Keq)

1.10V = 8.314J/K.mol × 298K/(2 × 96485C/mol) × ln(Keq)

ln(Keq) = 1.10×96485×2/(298×8.314) = 85.6754

Keq = e85.6754 = 1.6157×1037

= 1.62×1037 (Answer)